Sputnik The first artificial satellite to orbit the Earth was Sputnik I, launched October 4, 1957. The mass of Sputnik I was 83.5 kg, and its distances from the center of the Earth at apogee and perigee were 7330 km and 6610 km, respectively.

Find the difference in gravitational potential energy for Sputnik I as it moved from apogee to perigee. (Joules)
Answer #1

Potential energy is given by:

U = -GMm/r

at apogee, r = 7.33*10^6 m

at perigee, r = 6.61*10^6 m

dU = -GMm/rp - (-GMm/ra)

dU = GMm*(1/ra - 1/rp)

Using given values

dU = 6.67*10^-11*5.98*10^24*83.5*(1/(7.33*10^6) - 1/(6.61*10^6))

dU = -4.949*10^8 J